. A college drop out owes $42,000. Her loan has an interest rate of 6% a year and a ten year maturity. Determine her monthly payment. She currently has a job that pays $10 an hour for 160 hours a month. What percentage of her monthly before tax pay goes to paying off her loan? Use the table below for the following part: What percentage of her after tax income per year goes to paying her total loan payments for one year?

Single Filing Status

Income between

Marginal tax rate

$0 to $8,700

10%

$8,700 to $35,350

15%

$35,350 to $85,650

25%
Answer #1

1: Using financial calculator
Input:

PV = -42000,I/Y=6/12

N=10*12=120

Solve for PMT as 466.29

Monthly payment = $466.29

2: Monthly pre tax payment = $10*160 = 1600

Portion of payment going towards loan = 466.29/1600 = 29.14%

3: Annual income = 1600*12 = 19200

Tax = 15%

After tax income = 19200*(1-15%) = 16320

Portion of payment going towards loan = 466.29*12/16320

=34.29%